OpenStudy (anonymous):

what is the solution of the system of equations graphed below y=4x+11 y=-2x-1

OpenStudy (anonymous):

im trying to post the graph but it wont let me -.-

OpenStudy (anonymous):

y = 4x + 11 sub 4x + 11 in for y in the other equation y = -2x - 1 4x + 11 = -2x - 1 (add 2x to both sides) 4x + 2x + 11 = -1 (subtract 11 from both sides) 4x + 2x = -1 - 11 (combine like terms) 6x = -12 (divide by 6) x = -2 Now sub -2 in for x in either of the equations y = 4x + 11 y = 4(-2) + 11 y = -8 + 11 y = 3 check.. y = 4x + 11 3 = 4(-2) + 11 3 = -8 + 11 3 = 3 (correct) x = -2 and y = 3 or (-2,3)

OpenStudy (anonymous):

thank you (: